OpenStudy (anonymous):

Use substitution / find the ordered pair: x+2y=12 -x=-y-6

OpenStudy (amistre64):

1 2 = 12 -1 1 = -6 y = (12-6)/(1+2) 2 1 = 12 1 -1 = -6 x = (-12-12)/(-2-1)

OpenStudy (amistre64):

if its right, we get: y = 2 , and x = 8

OpenStudy (anonymous):

substitution: make the -x=...... positive by multiplying by -1 so x=y+6 substitute: (y+6)+2y=12 3y+6=12 3y=6 y=2 plug that in x=(2) +6 x=8 ordered pair: (8,2)

OpenStudy (amistre64):

for the record, the ordered pair really doesnt care how its found :)

OpenStudy (anonymous):

well he said use substitution :p

OpenStudy (amistre64):

yep, im just a dissenter for no apparent reason; kinda like that jimmy dean kid lol